OpenStudy (anonymous):

Can someone help me with this problem? f(x) = (3x - 2) / (4x +1) g(x) = (x+1) / (5x - 3) Value of (f+ g)(1)

OpenStudy (anonymous):

find \[f(1)\] and \[g(1)\] and add them up. can you find those numbers?

OpenStudy (anonymous):

Ohh is that all you do?

hero (hero):

Yep

OpenStudy (anonymous):

so i multiply it all out then add like terms?

hero (hero):

and also: (f+g)(x) = f(x) + g(x) so.... (f+g)(1) = f(1) + g(1)

OpenStudy (anonymous):

Alright thnx guys

hero (hero):

There are a few ways you can go about it. You can do it the easy way, or you can be a show off and do it the hard way

OpenStudy (anonymous):

alright well im gonna try and see how far i can get and come bak if i need to i guess

hero (hero):

Good Luck

hero (hero):

When is satellite going to pass level 100?

OpenStudy (anonymous):

just as a suggestion i believe in this case it would be much, much, much easier to evaluate each separately rather than doing a raft of annoying algebra and then evaluating

OpenStudy (anonymous):

never. i think you need \[10^{12}\] xp to get there

hero (hero):

What? Why?

hero (hero):

No one is going to catch you on medals until you're dead Satellite.

OpenStudy (anonymous):

71180/-565240 XP

OpenStudy (anonymous):

So for f(1) i got 1/5 and for g(1) i got 1. am i doin ok so far?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so its gonna end up being 6/5?

hero (hero):

So you did it the easy way. Now do it the hard way to check your answer.

OpenStudy (anonymous):

yup easy enuf right?

OpenStudy (anonymous):

yeah thnx every1=]